PITMinutes 08-13-2018

From CASA Wiki
Jump to: navigation, search


R. Bachimanchi, J. Benesch, M. Drury, A. Freyberger, R. Geng, C. Hovater, A. Hutton, G. Krafft, A. Kimber, C. Mounts, R. Nelson, T. Plawski, T. Powers, C. Reece, R. Rimmer, A. Seryi, and A. Solopova.

Agenda and Notes

1. Operations Status and Summary/Review (Freyberger)

During SAD two FEL modules were moved into NL07 and NL23 and one C100 was removed. In item [3] below Anna will tell us how they've started out.

Have started to turn everything on. Injector checked out to 100 muA but still adjustment to do. Beam has been through both linacs now. C100s seem to be generating vacuum excursions when operated at max gradient. The Gsets are being reduced on some cavities to as low as 8 MV/m to get stable ops. Not clear where we'll end up in energy yet.

WIld excursions observed in the JT valves for 1L21 and 1L25. Q0s blown?

LEMed north to 1050? Indicated trip rate of 5/hr. When gradients applied got vacuum excursions.

RF Heat and Q0 balancing are probably not in balance.

2. TTC Meeting and new info on particulates (Geng)

TTC Program

3. CEBAF Turnon Status (Solopova)

C100 Status at Turnon

4. C100 Refurbishment Planning (Krafft, Reilly)

C100 System Improvement Discussion

Items Request Email

C100 Refurbishment Brainstorming List

Tony's C100 Refurbishment Brainstorming ppt

Action Items

Leadership of quenches. Offline discussion.

C100 Refurbishment Project for the future?